### v2

# Finding Shared Decodable Concepts and their Negations in the Brain

Abstract
Prior work has offered evidence for functional localization in the brain; different anatomical regions preferentially activate for certain types of visual input. For example, the fusiform face area preferentially activates for visual stimuli that include a face. However, the spectrum of visual semantics is extensive, and only a few semantically-tuned patches of cortex have so far been identified in the human brain. Using a multimodal (natural language and image) neural network architecture (CLIP, Radford et al. (2021), we train a highly accurate contrastive model that maps brain responses during naturalistic image viewing to CLIP embeddings. We then use a novel adaptation of the DBSCAN clustering algorithm to cluster the parameters of these participant-specific contrastive models. This reveals what we call Shared Decodable Concepts (SDCs): clusters in CLIP space that are decodable from common sets of voxels across multiple participants.
Examining the images most and least associated with each SDC cluster gives us additional insight into the semantic properties of each SDC. We note SDCs for previously reported visual features (e.g. orientation tuning in early visual cortex) as well as visual semantic concepts such as faces, places and bodies. In cases where our method finds multiple clusters for a visuo-semantic concept, the least associated images allow us to dissociate between confounding factors. For example, we discovered two clusters of food images, one driven by color, the other by shape. We also uncover previously unreported areas with visuo-semantic sensitivity such as regions of extrastriate body area (EBA) tuned for legs/hands and sensitivity to numerosity in right intraparietal sulcus, sensitivity associated with visual perspective (close/far) and more. Thus, our contrastive-learning methodology better characterizes new and existing visuo-semantic representations in the brain by leveraging multimodal neural network representations and a novel adaptation of clustering algorithms.
1 Introduction
The study of the visual system has revealed a significant amount of knowledge about how the brain processes and organizes visual information. The early visual system has been well characterized for its ability to extract low level visual features (e.g. lines of different orientations). However, the picture becomes more murky as one moves beyond the early visual system. There is evidence for functional localization wherein large patches of cortex activate for specific visuo-semantic properties: areas of the brain preferentially activate for visual stimuli containing faces (Kanwisher et al., 1997) and places (Epstein and Kanwisher, 1998). There have been multiple recent attempts to identify new functionally localized regions of cortex with some converging evidence; for example, a food area near PPA has been identified (Khosla et al., 2022). So, while there has been progress towards parcelating the functional regions of the brain, only a few such broad visuo-semantic categories have been identified and they typically cover fairly large areas of cortex. This leaves open the possibility that we may be able to detect new visuo-semantic categories in the brain or refine existing categories, and that these categories or sub-categories may become more apparent with a data-driven approach that can combine data across multiple participants.
But combining brain responses across participants is not simple. Prior research typically focused on mapping to a shared latent space (e.g. during stimulus reconstruction) for which spatial localization in the brain is not required. In cases where spatial information is required, analyses are typically applied to data which has been warped to a common template brain, often losing the connection to the unique structure of each participant’s cortical layout. We present a novel method that finds semantically meaningful shared voxel clusters in each participant’s native brain space, allowing us to sidestep the voxel alignment problem. Derived voxel clusters can be mapped to a shared space for visualization after identification of shared semantic clusters.
Our data-driven analysis of fMRI recordings during image-viewing allows not only for identification of the images that most drive neural activity in a given area, but also of images which are most associated with below-average BOLD response. This provides stronger evidence for the possible function of cortical areas, in some cases refining the hypotheses one might draw from a set of images. Brain areas showing suppressed activity in response to a stimuli is consistent with recently reported “offsembles” of neurons (Pérez-Ortega_Akrouh_Yuste_2024) which show inhibited activity in response to e.g. visual stimuli, implying that, when interpreting neural representations, suppression of firing may be as important as firing.
We identify semantic concepts that drive activity in patches of cortex across multiple participants and find that their localization is fairly consistent across participants. To accomplish this, we introduce a novel variant of DBSCAN that combines data from multiple fMRI participants, identifying the shared decodable concepts (SDC) by specifically searching for SDCs present in multiple subjects.
2 Decoding CLIP-Space from Brain Images
To identify shared decodable concepts (SDCs) in the brain, we derived a mapping from anatomical brain space to a representational space in which we can explore semantic sensitivity to visual input. In this section we describe the components of this mapping: a dataset of fMRI recordings (NSD), a multimodal image-text embedding model (CLIP), and our method to map from per-image brain responses to their associated multimodal embeddings. We consider two models in this section, and verify that our proposed contrastive decoder outperforms a baseline ridge regression model.
2.1 fMRI Data
The natural scenes dataset (NSD) is a massive fMRI dataset acquired to study the underpinnings of natural human vision (Allen et al., 2022). Eight participants were presented with 30,000 images (10,000 unique images over 3 repetitions) from the Common Objects in Context (COCO) naturalistic image dataset (Lin et al., 2014). A set of 1,000 shared images were shown to all participants, while the other 9,000 images were unique to each participant. Single-trial beta weights were derived from the fMRI time series using the GLMSingle toolbox (Prince et al., 2022). This method fits numerous haemodynamic response functions (HRFs) to each voxel, as well as an optimized denoising technique and voxelwise fractional ridge regression, specifically for single-trial fMRI acquisition paradigms. Some participants did not complete all fMRI recording sessions and three sessions were held out by the NSD team for the Algonauts challenge. Further details can be found in Allen et al. (2022).
2.2 Representational Space for Visual Stimuli
To generate representations for each stimulus image, we use CLIP (Radford et al., 2021), a model trained on over 400 million text-image pairs with a contrastive language-image pretraining objective. CLIP consists of a text-encoder and image-encoder that jointly learns a shared low-dimensional space trained to maximize the cosine similarity of corresponding text and image embeddings. We use the 32-bit Transformer model (ViT-B/32) implementation of CLIP to create a 512-dimensional representation for each stimulus image used in the NSD experiment. We train a decoder to map from fMRI responses during image viewing to the associated CLIP vector for that same image. Notably, because CLIP is a joint image-language model, these CLIP vectors also correspond to text captions in the pretraining stage, which can be used to describe the images presented to the participants.
2.3 Data Preparation
Data Split
We split the per-image brain responses $\bm{X}$ and CLIP embeddings $\bm{Y}$ into training $(\bm{X}_{\text{Train}},\bm{Y}_{\text{Train}})$, validation $(\bm{X}_{\text{Val}},\bm{Y}_{\text{Val}})$, and test $(\bm{X}_{\text{Test}},\bm{Y}_{\text{Test}})$ folds for each of the 8 NSD participants. For each participant, the validation and test folds were chosen to have exactly 1,000 images with three presentations.  [^1] Some participants in the NSD did not complete all scanning sessions and only viewed certain images once or twice. These images are assigned to the training set, which varies in size across participants. Of the shared 1,000 images that all participants saw, 413 images were shown three times to every participant across the sessions released by NSD. These 413 images appear in each participant’s testing fold.
Voxel Selection
The NSD fMRI data comes with voxelwise noise ceiling (NC) estimates that can be used to identify reliable voxels. However, the NC is calculated using the full dataset. Therefore, these NC estimates should not be used to extract a subset of voxels for decoding analyses because they are calculated using images in the test set, which is a form of double-dipping (Kriegeskorte et al., 2009). We therefore re-calculated the per-voxel noise ceiling estimates specifically on our designated training data only. We selected voxels with noise ceiling estimates above $8\%$ variance explainable to use as inputs to the decoding model. This resulted in 6k-19k voxel subsets per participant (see supplementary section 6.1 for exact per-participant voxel dimensions input into the decoder) which were normalized according to section 6.5. In our visualizations, regions of the flattened brain surface in black represent voxels that have passed this voxel selection threshold.
2.4 Decoding Methodology
Decoding Model
The decoding model $\bm{g}(\bm{X};\bm{\theta})=\bm{\hat{Y}}$ is a linear model trained to map brain responses $\bm{X}=[\bm{x}_{1},\ldots,\bm{x}_{n}],\bm{x}_{i}\in\mathbb{R}^{v}$ to brain embeddings $\bm{\hat{Y}}=[\bm{y}_{1},\ldots,\bm{y}_{n}],\bm{y}_{i}\in\mathbb{R}^{512}$. Here $n$ is the number of training instances, and $v$ is the participant-wise number of voxels that pass the noise ceiling threshold. We optimize the brain decoder using the InfoNCE definition of contrastive loss (van den Oord et al., 2018), which is defined below in Equations 1 and 2.
$\text{Contrast}(\bm{A},\bm{B})=-\frac{1}{M}\sum_{i=1}^{M}log\bigg{(}\frac{exp(\bm{a}_{i}\cdot\bm{b}_{i}/\tau)}{\sum_{j=1}^{M}exp(\bm{a}_{i}\cdot\bm{b}_{j}/\tau)}\bigg{)}$ (1)
$\mathcal{L}_{\text{InfoNCE}}(\bm{A},\bm{B})=\frac{1}{2}[\text{Contrast}(\bm{A},\bm{B})+\text{Contrast}(\bm{B},\bm{A})]$ (2)
In this definition $\bm{A}=[\bm{a}_{1},\ldots,\bm{a}_{n}]$ and $\bm{B}=[\bm{b}_{1},\ldots,\bm{b}_{n}]$ are embeddings for two modalities representing the same data points, the $\cdot$ operator represents cosine similarity, and $\tau$ is a temperature hyper-parameter. The loss is minimized when the distance between matching embeddings is small, and the distance between mismatched embeddings is large. In the original CLIP setting, $\bm{A}$ and $\bm{B}$ represent embeddings for images and corresponding text captions. In our implementation, we apply the contrastive loss to image embeddings computed by a pretrained frozen CLIP model and CLIP embeddings predicted from fMRI, i.e. we optimize $\min_{\bm{\theta}}\>\mathcal{L}(\bm{\hat{Y}},\bm{Y}_{\text{CLIP}})=\mathcal{L}(g(\bm{X};\bm{\theta}),\bm{Y})$. An illustration of the decoding procedure is given in Figure 11 (additional details on the decoding procedure are given in Section 6.1).
Evaluation
We evaluate our models using top-k accuracy. Figure 12 shows the results of this evaluation. The contrastive decoder outperforms ridge regression across all values of $k$. Recall that our end goal is to identify shared decodable concepts (SDC) in the brain. Our methodology for this task relies on the predicted CLIP vectors, and so our subsequent analyses require an accurate decoding model.
3 Finding Shared Brain-Decodable Concepts and their Representative Images
In this section, we describe our methods for finding and interpreting areas of the brain that respond to semantically similar images, and that are consistent across participants. We approach this problem by analyzing the weight matrices $\bm{W}^{(k)}$ from the optimized linear decoders described in section 3.4 $g^{(k)}(\bm{x}^{(k)})=\bm{W}^{(k)}\bm{x}^{(k)}=\bm{y}$ for all subjects $k\in\{1\ldots 8\}$. First, we observe that the decoder’s linear transformation $\bm{W}^{(k)}\bm{x}^{(k)}=\sum_{i=1}^{v}\bm{w}_{i}^{(k)}\cdot{x_{i}}^{(k)}$ is simply a summation of parameter vectors $\bm{w}_{i}^{(k)}\in\mathbb{R}^{512}$ that are scaled by brain response values $x_{i}^{(k)}\in\mathbb{R}$. This means that when $x_{i}^{(k)}$ is above or below baseline activation, the CLIP dimension represented by $\bm{w}_{i}^{(k)}$ is increased or decreased in the brain-decoded embedding. This motivates the key to our analysis: we view the parameter vector $\bm{w}_{i}^{(k)}$ as a CLIP vector that represents a brain-decodable concept for voxel $i$. This allows us to use cosine distance $d(\bm{w}_{i}^{(k)},\bm{w}_{j}^{(r)})$ between the weight vectors for voxels $i,j$ from participants $k,r$ as a measure of similarity between the decodable concepts of brain voxels across participants. Our objective is to apply a clustering algorithm using this metric in order to find areas in the brains of multiple participants that have a high conceptual similarity. We refer to these underlying concepts as shared decodable concepts (SDCs). We interpret the SDCs by retrieving a set of representative images associated with the brain-decoded CLIP vectors that are closest to the centroid of each SDC cluster. A schematic of the algorithm we apply is given in Figure 1.
Figure 1: Deriving SDC clusters. (a): The participant-specific linear decoders derived in Figure 6.1. (b): Our modified DBSCAN clustering procedure is applied to the linear decoders (See Figure 2 for details). (c): Our DBSCAN procedure derives binary masks over the voxels in the linear decoders for a specified number of clusters (of which one is highlighted in orange) (d): The rows corresponding to the selected voxels in the binary masks are extracted from the linear decoder matrices. (e): The 512-dimensional representations from the previous step are averaged over voxels and participants to derive a cluster centroid for each cluster derived from DBSCAN. We visualize the cluster centroid for the first DBSCAN cluster. (f): The linear decoders from (a) are applied to held-out fMRI data per-participant. Each participant saw images multiple times, so the matrix of predicted CLIP embeddings $\hat{Y}^{k}_{Test}$ is averaged over these repetitions and all linear decoding matrices are stacked (across participants) to give $\hat{Y}^{AVG}_{Test}$. (g): Cosine distance is calculated between the cluster centroids (e) and the brain-derived CLIP embeddings (f). (h): The images most associated with the cluster centroids (positive images) and most negatively associated with the cluster centroids (negative images) are identified. Positive / negative images for the SDC cluster pictured here appears to correspond to global vertical/horizontal orientation in the associated images. (i): Color-coded participant-specific voxel clusters are displayed on a flatmap of the brain’s cortical surface in common fsaverage space (overlapping areas are displayed in white). Regions of interest labels are highlighted on the flatmap image in white outlines. For the specified cluster (e), whose positive / negative images are associated with orientation, the flatmap indicates bilateral shared voxel clusters in early visual cortex.
3.1 Cross Participant Clustering
To discover shared brain-decodable concepts we apply a novel clustering method to the per-voxel model parameter vectors $\bm{w}_{i}^{(k)}$ across all participants. We base our clustering method on the density-based spatial clustering of applications with noise (DBSCAN) (Ester et al., 1996) algorithm, which is parameterized by a neighborhood size $\varepsilon$ and a point threshold minPts. The algorithm can be summarized using the following steps: (1) Points that have at least minPts points in their $\varepsilon$-neighborhood are marked as core points; (2) construct a graph $\mathcal{G}$ where the vertices are core points, and there are edges between core points that have a distance less than $\varepsilon$; (3) clusters are formed by finding the connected components of $\mathcal{G}$ and (4) the remaining non-core points are added to clusters if they are within the $\varepsilon$-neighborhood of a core point, otherwise, they are marked as outliers not belonging to any cluster.
In our modified cross-participant DBSCAN, we redefine the core point threshold minPts as minNeighbors. Then, in step (1) a point $\bm{w}_{i}^{(k)}$ is a core point if there are points from at least minNeighbors other participants within its $\varepsilon$ neighborhood. With this modification, a core point now identifies voxels that represent a brain-decodable concept that is shared with at least minNeighbors participants. Since there are 8 participants in the study, the maximum value of minNeighbors is 7. Steps (2) to (4) proceed as above.
Our second modification is to apply a within-participant expansion of clusters. We define a new hyperparameter $\varepsilon_{\text{expansion}}$ and apply a final expansion step: (5) All points inside the $\varepsilon_{\text{expansion}}$ neighborhood of a point in a cluster become members of that cluster with the constraint that they must be from the same participant. The addition of this step was motivated by our early experiments, where we noticed that there were sometimes only one or two voxels belonging to certain participants for each cluster. This allows the cluster boundaries to grow slightly larger for each participant, and we found that this did not significantly change the cluster location or semantic interpretations. Figure 2 outlines the application of our modified DBSCAN algorithm to modeling fMRI data.
We present results using a fixed value for $\texttt{minNeighbors}=3$. This means that each cluster must include at least 4 out of 8 participants in the study. This allows for the discovery of SDCs that may show up less reliably in some participants. Furthermore, we use a range of values for $\varepsilon\in\{0.5,0.55,0.6,0.65\}$ to explore clusters at different density scales. We set the expansion neighborhood size $\varepsilon_{\text{expansion}}=\text{min}(\varepsilon+0.05,0.65)$ so that the cluster boundaries can grow slightly larger than the baseline neighborhood size $\varepsilon$, but not exceeding 0.65 as we noticed neighborhoods become over-connected as $\varepsilon$ approaches 0.7. The number of cross-participant clusters found was 9, 14, 12, 11 for $\varepsilon$ values of 0.5, 0.55, 0.6, 0.65 respectively.
To reduce the impact of random initialization and combat the noise inherent in fMRI, we re-trained the decoder model 50 times for every participant and compute the average of the resulting parameter vectors. In other words, each averaged parameter vector $\bm{w}_{i}^{(k)}$ used in our analysis represents the average of 50 concepts that a voxel could represent depending on the random initialization of the decoder. We found that this significantly improved the quality and consistency of the clusters revealed by our modified DBSCAN algorithm.
Figure 2: Illustration of our DBSCAN variant applied to multi-participant fMRI data. There are 3 participants in this example and $\texttt{minNieghbors}=2$. (a): A zoomed-in view of a cluster with three core points. The outer ring around each point shows its $\varepsilon$-neighborhood and whether it is a core, border, or outlier point. Black arrows emphasize points that are neighbors. The points with neighboring points from at least 2 other distinct participants are marked as core points. Non-core points that neighbor core points are added to the cluster as border points. The remaining points are marked as outliers. (b): A zoomed-out sketch of a set of points that form 3 clusters. Since $\texttt{minNieghbors}=2$, a high-density region will form clusters if and only if it contains points from at least 3 participants.
3.2 Selecting Representative Images
To interpret the semantic meaning of an SDC cluster, we first computed the cluster centroid by taking the mean of all parameter vectors within the cluster $\bm{\bar{w}}=\frac{1}{|I|}\sum_{(i,k)\in I}\bm{w}_{i}^{(k)}$ where $I$ is a set of tuples identifying the voxel and subject indices for an SDC cluster. When the cluster voxels are above or below baseline activation, the CLIP dimension represented by $\bm{\bar{w}}$ is increased or decreased in the brain-decoded embedding, respectively. We note that DBSCAN has the ability to create non-globular clusters for which the centroid might not be a good representative. In our experience the centroid is within $\varepsilon$ of a core point within the cluster, implying that it is likely fairly representative of the cluster.
In order to find the representative images for an SDC cluster, we can consider the nearby image-embeddings to the cluster centroid $\bm{\bar{w}}$. Instead of using the original CLIP embeddings for each image, we instead used the subject-specific brain-decoded embeddings $\bm{\hat{Y}}_{\text{Test}}^{(k)}$. This allows us to focus specifically on the information that can be decoded from fMRI, which could be a subset of the information represented by CLIP space. For images that were viewed more than once we average the embeddings within and across participants. These averaged brain-decoded embeddings from held-out test data are denoted by $\bm{\hat{Y}}_{\text{Test}}^{\text{AVG}}$. To select nearby images we define $D(\bm{w},\bm{Y})=\{d(\bm{w},\bm{y})|\bm{y}\in\bm{Y}\}$ which is the set of distances between a CLIP vector $\bm{w}$ to a set of CLIP embeddings $\bm{Y}$. We take the cosine distances between the SDC cluster centroid $D(\bm{\bar{w}},\bm{\hat{Y}}_{\text{Test}}^{\text{AVG}})$ and the negated centroid $D(-\bm{\bar{w}},\bm{\hat{Y}}_{\text{Test}}^{\text{AVG}})$ and retrieve the images corresponding to the smallest values in these sets as the positive and negative representative images respectively.
4 Results
The clustering method described in chapter 3 not only identifies previously discovered functional areas but additionally identifies new ones, both of which we explore in this section. We emphasize that there is no constraint that voxels within a shared visual semantic cluster be spatially adjacent in the brain, yet during visualization we consistently find contiguous patches both within and across participants. We can localize these patches of voxels to specific regions of interest (ROIs) both functionally and anatomically (See Section 6.3 for further details).
4.1 Faces
One of the first reported functionally localized areas for higher-order vision was the fusiform face area (FFA)  (Kanwisher et al., 1997). Our method identifies a face-related concept (Figure 4.1 $\varepsilon=0.55$, cluster 7) that is localized to FFA and includes voxels from all 8 participants. This cluster also has some voxels in the extrastriate body area (EBA), likely because many face images include part or all of a person’s body. We note that the positive representative images are not exclusively human and include a range of animal faces. The images often depict people eating and handling food, or holding other objects such as toothbrushes or cell phones.
Interestingly, the negative representative images often display bodies, but there is a striking lack of clearly visible faces. There are several examples where people are visible, but their faces are obscured or they are facing away from the camera. This suggests that there may be a dip in FFA activity for images where a person’s face might be expected (and thus FFA is primed to activate) but not clearly visible.  [^2] Thus the brain’s representation for the opposite of a face is a scene with the conspicuous absence of faces.
Figure 3: Cluster 7 ($\varepsilon=0.55$). Positive images are strongly associated with faces, while the negative images represent depictions of people whose faces are not visible. Voxel clusters are primarily found in bilateral FFA and EBA.
Figure 4: Cluster 0 ($\varepsilon=0.55$). Positive images are associated with food and color. Negative images are entirely grayscale. Voxel clusters span bilateral FFA, V4, and PPA.
4.2 Food and Color
Previous work has specifically noted the correlation of very colorful images with food-related images, and attempted to define food areas in the absence of color  (Jain et al., 2023). Our method also identifies a large possibly food-related cluster that spans FFA, PPA, and V4 (cluster 0, $\varepsilon=0.55$, Figure 4). At first inspection most of the positive representative images are food-related. However, we also observed many vibrant and colorful positive images that contained no food. Strikingly, the negative representative images are entirely gray-scale, providing very strong evidence that this cluster may be related to color. Thus, we speculate that this cluster is not specifically related to the identification of food, but might rather correspond to any colorful image. We discuss two other food-related clusters in supplementary section 6.8.
4.3 Bodies
Our method identifies five notable body-related areas at $\varepsilon=0.55$ in and around EBA (downing_2001). The positive representative images for cluster 2 (Figure 5) show people and animals outside with an emphasis on legs and active movement. The negative images typically depict people indoors sitting with their legs obscured by tables. Cluster 11 (Figure 6) has a similar emphasis on hands instead of legs. People are displayed in a variety of contexts with their hands clearly visible, while the negative images are exclusively non-primate animals without hands. This cluster also shows indoor/outdoor contrast in the representative image groups and so there is some PPA activation. In Section 6.7 we outline further body-related shared voxel clusters associated with motion and a distinction between single individuals and groups of people in a sub-region of EBA.
Figure 5: Cluster 2 ($\varepsilon=0.55$). Positive images are strongly associated with presence of legs, while negative images are typically people at tables whose legs are obscured. Voxel clusters are primarily in bilateral EBA.
Figure 6: Cluster 11 ($\varepsilon=0.55$). Positive images are strongly associated with hands and hand motion, while negative images are associated with animals (no hands). Voxel clusters are primarily in bilateral EBA and FFA.
4.4 Horizontal/Vertical Cluster
Very early work on the visual system discovered the tuning of the early visual system for lines of a particular orientation. Two clusters emerge that reflect this tuning, both at $\varepsilon=0.65$. Cluster 1 (Figure 7) has a strong horizontal component with strong horizon lines or large objects spanning the middle of the visual field creating a horizon-like line. Notably, the negative images are images with a strong vertical component. The inverse is true for cluster 7 (Figure 8) which shows strong verticality in the positive images, and horizontal in the negative. Interestingly some of the negative images from cluster 1 appear as positive images in cluster 7, and vice versa. The localization of both clusters is V2/V3.
Figure 7: Cluster 1 ($\varepsilon=0.65$). Positive images are associated with a horizontal mid-line, while negative images display a vertical mid-line. Voxel clusters are primarily in right hemisphere V2.
Figure 8: Cluster 7 ($\varepsilon=0.65$). Positive images are associated with a vertical mid-line, while negative images display a horizontal mid-line. Voxel clusters are primarily in bilateral V3.
4.5 Repeated Elements (Numerosity) Cluster
We observed that cluster 12 ($\varepsilon=0.55$, Figure 9) has positively associated images typically with repeated elements of related items, while negative images typically contained singular instances of items. We suggest that this SDC could be related to quantity processing and numerosity. We observed voxel clusters primarily in right OPA and IPS regions. As with other clusters, naturalistic images often implicates functional areas associated with place (OPA). The IPS has been widely reported in prior work to be associated with quantity processing, e.g. grammatical number processing Carreiras et al. (2010), mathematical processing deficits Ganor-Stern et al. (2020) and numerical processing Koch et al. (2023)). Our observations in the visual domain provide converging evidence that the right IPS is associated with quantity processing across multiple modalities.
Figure 9: Cluster 12 ($\varepsilon=0.55$). Positive images are associated with repeated elements, while negative images are single objects or individuals. Voxel clusters are primarily located in OPA and IPS.
Figure 10: Cluster 2 ($\varepsilon=0.65$). Positive images display close-ups of large animals or objects. Negative image show objects at a distance. Voxel clusters are located between EBA and V4.
4.6 Close vs Far Cluster
With $\varepsilon=0.65$, cluster 2 (Figure 10) seems to represent big things (airplanes, busses) and close up pictures of larger animals (zebras, elephants). The negatively associated images are scenes that usually depict significant depth, including paths leading into the hills with animals or people in the distance. Sarch et al. (2023) explored the representation of image depth in cortex and also found that similar areas of cortex respond strongly to images with objects very close to the camera. In addition, Luo et al. (2023) reported that this general area is sensitive to relatively large objects.
5 Conclusion
We proposed a novel mechanism to identify shared decodable concept clusters in a large multi-participant fMRI dataset of natural image viewing (Allen et al., 2022). To achieve this, we applied a contrastive learning approach using multimodal embeddings from CLIP (Radford et al., 2021) and a novel adaptation of the DBSCAN clustering algorithm (Ester et al., 1996). Our method identifies previously reported functional areas that code for various visuo-semantic categories, as well as identifying novel shared categories. Furthermore, our method exposes interpretability by identifying the most / least associated images with CLIP-based cluster centroids. In the case of faces, we find a set of positive / negative images where the negative images depict scenarios where a face might be expected, but is not visible. This suggests that the brain responses for this SDC cluster are organized in a way that specifically codes for the presence of faces in expected visual depictions. Food is a concept that is highly confounded with other visual criteria such as shape and color. By providing the most negatively associated images for a SDC cluster, we enabled easy interpretation of semantic coding, e.g. in the case where the negated images to this SDC cluster are all grayscale images (see Section 4.2). We further identify SDC clusters specific to bodies in motion, in groups, as well as hands and legs. We identify SDC clusters related to repeated elements (numerosity), sensitivity to indoor / outdoor scenes, cardinal orientation of objects in images (horizontal / vertical) and gradations of shape information (soft vs hard) and visual perspective (close vs far). Our method showcases how various techniques from machine learning can be employed to better characterize visual representations of semantic selectivity in the human brain.
5.1 Broader Impacts
Methods that decode neural activity could have substantial impacts on both neuroscience and broader society. Identifying new category-selective brain regions could assist in diagnosing neurological and psychiatric conditions, surgical planning, development of brain-computer interfaces, and potentially helping individuals with locked-in syndrome or related disabilities to better communicate.
However, there are also significant ethical considerations with technologies that decode neural responses. If these techniques are used outside controlled research environments, there is a risk that they could be used for intrusive monitoring of mental states. As with all emerging technologies, responsible deployment is needed in order for society as a whole to maximize benefit while mitigating risks.
5.2 Limitations and Future Work
While our approach offers many advantages, it is not without drawbacks. Firstly, we are limited by the stimulus images that were chosen for the NSD experiment. Although the use of a contrastive loss function helps mitigate this, our model may still be biased toward over-represented categories in the stimulus set. Additionally, the use of CLIP may introduce its own biases, as some signal in the brain responses may not be fully captured by the CLIP embeddings. The use of fMRI also presents certain limitations. Despite its high spatial resolution, its limited temporal resolution restricts the ability to capture rapid neural activity that may play a crucial role in visual processing. Future work could address these limitations by utilizing larger and more diverse fMRI datasets, incorporating higher temporal imaging techniques such as electroencephalogram (EEG) or electrocorticography (ECoG), and exploring alternative stimulus representations beyond CLIP.
Additionally, our clustering approach might merge concepts within regions of CLIP space that have relatively uniform densities, potentially missing interesting SDCs. This issue could be addressed by adapting the heirarichcal version of DBSCAN, or by using different criteria to merge the cross-participant core points into clusters.
Furthermore, we acknowledge that our method is not a replacement for traditional hypothesis-driven experiments. It will be important for future work to test for alternative hypotheses for what might be driving responses in SDC clusters. Another interesting research direction is the application of this method to other sensory modalities beyond vision, such as auditory or somatosensory stimuli.